Determine the general solution by solving the ODE system

2r+y

= 4y
Answer #1

Given tat dn d t んてニ + Compare wrth dy By Au and lay equation uiven (artbm 9ib2-926o m b1 e2 : -, b2 = q 1Heme a 2 m2 (2+4) ma solutton o we ofeourre A =118 0 thae St nこe duia solution りon Seek a Seconeel h T Solutio af Now Lee Corm (A+ A2)e =(B4 B2(Aa+4B2]t+Alt Aet 401にC A2-B2)t Ai-B1t B2 o Compae tee Coetd af 4 A +A2T 4Bj - OG) Ai-6」+B2こ0 A2-B2 O A2ニ6と Pet in 5 B2- 0 2